# Equations of State, Thermodynamics, and Miscibility Curves for Jovian Planet and Giant Exoplanet Evolutionary Models

### v1

@@ Abstract
The equation of state of hydrogen-helium (H-He) mixtures plays a vital role in the evolution and structure of gas giant planets and exoplanets. Recent equations of state that account for hydrogen-helium interactions, coupled with hydrogen-helium immiscibility curves, can now produce more physical evolutionary models, such as accounting for helium rain with greater fidelity than in the past. In this work, we present a set of tools for planetary evolution111Available at https://github.com/Rob685/hhe_eos_misc that provides a Python interface for tables of useful thermodynamic quantities, state-of-the-art H-He equations of state, and pressure-dependent immiscibility curves. In particular, for a collection of independent variable choices, we provide scripts to calculate a variety of thermodynamic derivatives used to model convection and energy transport. This centralized resource is meant to facilitate and consolidate giant planet structural and evolutionary modeling going forward.
@@ 1 Introduction
The structure and evolution of gas giant planets and brown dwarfs depend on the properties of hydrogen and helium at high pressures (Hubbard, 1970; Hubbard & Slattery, 1971; Hubbard & Dewitt, 1985; Burrows et al., 1997, 2001; Fortney & Hubbard, 2003, 2004; Militzer et al., 2008; Nettelmann et al., 2008, 2012, 2015; Hubbard & Militzer, 2016; Mankovich et al., 2016; Püstow et al., 2016; Miguel et al., 2016; Debras & Chabrier, 2019; Mankovich & Fortney, 2020; Debras et al., 2021a; Mankovich & Fuller, 2021; Howard et al., 2023), as well as their atmospheres (Burrows et al., 1997; Fortney et al., 2011; Burrows, 2014; Chen et al., 2023). The equation of state (EOS) most employed in the past was published by Saumon et al. (1995, SCvH95). This EOS incorporated the so-called “chemical picture” of the constituent species, obtained by minimizing the Helmholtz free energy of its constituents. This model estimated a gradual disassociation and ionization of hydrogen with increasing pressure, thus accounting for the increasing densities and even the predicted hydrogen first-order phase transition within Jupiter. Subsequent ab initio calculations by Militzer & Hubbard (2013, MH13) and Nettelmann et al. (2012), which included non-ideal interaction effects of hydrogen-helium mixtures and partial ionizations, were an important step forward. Models using updated EOSes by MH13 and Nettelmann et al. (2012) and the SCvH95 EOS resulted in significant differences in the interior structures of Jupiter and Saturn (see Miguel et al., 2016; Militzer & Hubbard, 2013; Hubbard & Militzer, 2016; Miguel et al., 2022). Further complications arose when applying such EOSes to such modeling due to 1) the phase separation of helium related to the metallization of hydrogen at high pressures; 2) the presence of a higher-Z core; and 3) the presence of a component of heavy elements in the gaseous mantle. The distribution of the latter is the subject of intense debate following the recent interpretations (Leconte & Chabrier, 2012, 2013; Vazan et al., 2016; Wahl et al., 2017; Folkner et al., 2017; Vazan et al., 2018; Stevenson, 2020; Helled et al., 2022; Militzer et al., 2022; Howard et al., 2023) of the Juno (Bolton et al., 2017a, b) gravity data. To this end, two primary areas of focus are an improved H-He EOS and a more accurate H-He phase separation curve.
@@
The first area, the H-He EOS, demands ab initio calculations of relevant thermodynamic quantities. The computational costs have impeded the creation of a comprehensive tabular EOS for a wide range of hydrogen-helium mixtures, temperatures, and pressures. As a result, current EOSes employed in structural and evolutionary codes, such as those of SCvH95, (Nettelmann et al., 2012; Becker et al., 2013), Chabrier et al. (2019, CMS19) and Mazevet et al. (2022, MLS22), rely on idealized mixing laws to compute the entropy of H-He mixtures. However, a newer class of models includes non-ideal effects of the entropy of mixing. These terms have recently been estimated by Howard & Guillot (2023, HG23) based on the calculations of MH13, who calculated a H-He mixture EOS for a helium mass fraction () of 0.245. This resulted in estimates for the non-ideal corrections for any helium mass fraction for the CMS19 and MLS22 EOSes.
@@
Between the different EOS and phase separation calculations by various authors, planetary modelers have a wide range of options. In this paper, we collect many major EOS and miscibility databases required for giant planet evolutionary and structural calculations and describe their strengths and weaknesses222We omit the H-He EOS of Nettelmann et al. (2012) and Becker et al. (2014) due to the lack of entropy calculations available in their tables. An integration technique to obtain the entropy for these tables depends on choices of initial thermodynamic values, making the comparisons between EOSes uncertain and subsequent derivative calculations subject to this uncertainty.. We invert each EOS on a dense grid and calculate all the necessary quantities and derivatives that play a role in the physics, structure, thermal evolution, and convection criteria within Jupiter and Saturn in particular, and giant exoplanets in general. Each EOS is calculated along pressure-temperature, density-temperature, and entropy-density axes that span the thermodynamic range of the interiors of gas giants. As discussed in Section 5, a proper treatment of convection requires the calculation of various derivatives. In addition to the various EOS tables, we furnish useful tables of the hydrogen-helium immiscibility curves of Lorenzen et al. (2009, 2011) and Schöttler & Redmer (2018). In this way, our work provides a more centralized and standardized approach by collecting the most recent EOS and H-He immiscibility data in a self-consistent format.
This paper is organized as follows. Section 2 discusses the various hydrogen-helium (H-He) EOSes, compares their properties, and discusses their applications. In Section 3, we provide an overview of the available miscibility calculations in the literature, compare the different demixing temperature calculations of each of the sources, and discuss their differences vis à vis planetary evolution. In Section 4, we briefly discuss the thermodynamic consistency of the EOSes published by Saumon et al. (1995) and Chabrier et al. (2019). In Section 5, we discuss the various derivatives necessary for generalized convection and we review the different derivatives necessary for energy calculations in a planetary model. Finally, in Section 6, we provide a summary discussion and concluding remarks. In the appendices, we derive various of the relationships provided in the text.
@@ 2 Hydrogen and Helium Equations of State
Historically, the EOS of hydrogen and helium at high pressures has been divided into the “physical” picture and the “chemical” picture. In the physical picture, electrons and nuclei interact through Coulomb potentials, and the EOS is obtained by solving the resulting quantum many-body problem. In the chemical picture, atoms remain bound and interact through pair potentials. However, at ionization pressures and temperatures, electrons become unbound and delocalized. In the high temperature regime, electrons become unbound and the physical picture is tractable (Rogers, 1981; Rogers & Young, 1984), and interactions under the Thomas-Fermi approach (Myers & Swiatecki, 1990, 1996). However, in the low-temperature and high-pressure regime relevant to gas giant planets, the physical picture remained intractable due to partial ionization. The chemical picture presents a simplified model a physical model. In the past, the chemical picture model of SCvH95 was the most commonly used H-He EOS. However, with improvements in computation, rigorous physical (or “ab initio”) models have been published. Most notably, MH13, CMS19, Chabrier & Debras (2021, CD21), and MLS22 have published EOSes obtained from ab initio calculations. As discussed in the introduction, these previous EOSes handle arbitrary H-He mixtures via the ideal additive volume law (AVL; see Section 2.1). This approximation was improved by HG23, who have calculated nonideal interaction corrections based on the MH13 results for the CMS19 and MLS22 EOSes. In this section, we discuss the properties of the historical SCvH95, the ab initio MH13, and the most current CMS19 EOSes and compare their properties333The MLS22 EOS result is a hydrogen EOS table benchmarked with structural calculations of the interior of Jupiter inferred from the Juno data. Though we present this EOS and provide the same quantities and derivatives for it, we focus here on the CMS19 and SCvH95 EOSes..
@@ 2.1 Saumon, Chabrier, and van Horn (1995)
Saumon et al. (1995, SCvH95) employed the chemical picture and generated an EOS that covered a temperature range of and a density range of . They calculated pure hydrogen and helium tables and relied on the additive-volume law (AVL), also referred to as the “volume-addition law,” to handle a general helium mass fraction dependence. Intensive variables, such as temperature and pressure, must be uniform throughout a system at equilibrium. Extensive variables, such as volume and entropy, are additive for a system in equilibrium. Therefore, the AVL considers a combined system of equations of state and assumes they are in thermal and pressure equilibrium. AVL cannot account for interactions between hydrogen and helium, is not reliable in regions of partial ionization, and it does not account for phase separation (see Section 6 of SCvH95 for more detailed descriptions of its limitations). Despite these limitations, until recently, SCvH95 was the most used EOS in the study of low-mass stars, brown dwarfs, and giant planets (e.g., Burrows et al., 1995; Guillot et al., 1996; Saumon et al., 1996; Burrows et al., 1997; Chabrier & Baraffe, 1997; Baraffe et al., 1998; Burrows et al., 2001; Fortney & Hubbard, 2003, 2004; Fortney et al., 2011)
@@ 2.2 Militzer & Hubbard (2013)
The Militzer & Hubbard (2013, MH13) EOS spans a temperature range of K and a density range of g cm. As their model is a physical model, they avoid the limitations of the low-temperature and high-pressure regime inherent in the chemical model of SCvH95. They use Density Functional Theory Molecular Dynamics (DFT-MD), coupled with the thermodynamic integration (TDI) technique, to obtain the entropy directly through the difference in Helmholtz free energy between hydrogen and helium mixtures. To account for interactions between species and electrons, they used intramolecular pair potentials derived at a particular density and temperature and the Perdew-Burke-Ernzerhof (Perdew et al., 1996, PBE) exchange-correlation functional.
@@ 2.3 Chabrier, Mazevet, and Soubiran (2019)
Chabrier et al. (2019, CMS19) constructed pure hydrogen and helium EOS tables spanning wider density and temperature ranges, namely g cm and K. For densities lower than 0.05 g cm, they use the SCvH95 hydrogen EOS, the hydrogen EOS of Caillabet et al. (2011) for g cm, and the EOS of Chabrier & Potekhin (1998) for densities greater than 10 g cm.
As SCvH95, their EOS rely on the AVL for helium mixtures and include only the ideal entropy of mixing. CD21 attempted to incorporate the non-ideal interaction effects of MH13 into the CMS19 EOS using methods from Miguel et al. (2016). Using the AVL, they obtained an “effective” hydrogen EOS from MH13 using the CMS19 helium EOS to obtain a mixuture for . With this correction, the CD21 results agree at higher densities with MH13 (see Figure 9 of CD21).
@@
and requiring that these corrections agree with MH13 at and with CD at . The corrections, and , are assumed to be proportional to the products of hydrogen and helium mass fractions.
The CMS19 EOS with the HG23 corrections (CMS19+HG23) is the most current development of an H-He EOS that spans all relevant thermodynamic space, helium fractions, and nonideal interactions. Based on this, we advocate CMS19 + HG23 EOS as a current default baseline. In Figure 1, we compare the CMS19+HG13 and MLS22+HG23 EOSes with the SCvH95, MH13, CMS19, and CD21 tables, illustrating the differences among them. The top panel mirrors Figure 4 of HG23, comparing the EOSes at a helium fraction of and a specific entropy of baryon. The bottom panel shows the temperature differences due to the non-ideal interactions at different entropy values, as noted above, for each profile. The density deviations from SCvH95 increase with decreasing entropy, particularly between 0.5 and 7 Mbar. This pressure region is expected to exhibit H-He phase separation at a sufficiently low temperature (Stevenson, 1975; Stevenson & Salpeter, 1977). In the next section, we discuss the current understanding of H-He immiscibility in this context.
@@ 3 Hydrogen and Helium Immiscibility
Early evolutionary calculations for Saturn (Pollack & Graboske, 1976; Pollack et al., 1977) could not reproduce its current heat flux. To account for this apparent excess heat, Stevenson (1975) proposed the formation of He-rich droplets that rain into the interior (also see Stevenson & Salpeter, 1977). One of the first attempts to capture and apply the demixing of hydrogen and helium as a function of helium abundance was done by Hubbard & Dewitt (1985), who applied Monte Carlo simulations for a fully ionized hydrogen and helium plasma. The low-pressure regime was studied by Schouten et al. (1991) via Monte Carlo methods using pair potentials for interactions. Pfaffenzeller et al. (1995) calculated the demixing temperatures using molecular dynamics (MD) and found a demixing temperature-pressure dependence opposite that of Hubbard & Dewitt (1985). However, the demixing temperatures remained too cold for Saturn to ever reach the immiscibility regions (Pfaffenzeller et al., 1995; Fortney & Hubbard, 2003). Fortney & Hubbard (2003) created an ad hoc phase diagram, adjusting the demixing temperatures of Pfaffenzeller et al. (1995), to match the present effective temperature of Saturn. More recently, H-He phase diagrams have been calculated by Lorenzen et al. (2009, 2011, L0911), Morales et al. (2009, 2013), and Schöttler & Redmer (2018, SR18). These ab initio phase diagrams are discussed in the next sections.
@@
Lorenzen et al. (2009, L0911) conducted the first ab initio immiscibility gap model. They used DFT coupled with Car-Parrinello MD, which was also used by Pfaffenzeller et al. (1995), with a PBE exchange-correlation functional and pair potentials for hydrogen and helium. They calculated the demixing temperatures for pressures between 4 Mbar and 24 Mbar and expanded their model in Lorenzen et al. (2011) to include 1 and 2 Mbar. In all cases, they relied on linear mixing to calculate the enthalpy of mixing to then infer the change in the Gibbs free energy, . This was then minimized to constrain the H-He miscibility gaps. They linearized the change in the enthalpy with
@@
where ] is the change in the ideal entropy of mixing. This term does not account for any non-ideal interactions in their EOS.
@@
The phase diagrams from L0911 predict a large miscibility gap in Saturn (Püstow et al., 2016) and little to no gap in Jupiter (Mankovich et al., 2016) at the present age. Püstow et al. (2016) claim that the L0911 diagrams need corrections of K and K for Jupiter and Saturn, respectively. In the next section, we discuss the work of Morales et al. (2009, 2013) and their inclusion of non-ideal interactions in the phase diagram.
@@
Morales et al. (2009) constructed the first H-He phase diagram that included non-ideal interaction effects in the entropy of mixing, following a similar methodology to that of Militzer (2009) and MH13. Instead of the Car-Parrinello MD, Morales et al. (2009) used the Born-Oppenheimer MD coupled with the PBE exchange-correlation functional, and used TDI to infer the entropy using the Helmholtz free energy difference (also used by MH13). Morales et al. (2013) expanded their work to pressures below 0.4 Mbar, and claimed that the linear mixing assumption leads to significant disagreements with their direct calculation methods at high helium mass fractions and high pressures.
This work led to colder demixing temperatures compared to L0911. While Jupiter had a small miscibility gap between Mbar and Mbar with the L0911 phase diagram, the Morales demixing temperature results were too cold to allow Jupiter any miscibility gap (Morales et al., 2013). For consistency, Hubbard & Militzer (2016) used the Morales phase diagram, since both the EOS and the demixing temperatures were calculated in a similar fashion. Hubbard & Militzer (2016) then claimed that helium rain occurs within Jupiter at an entropy value of baryon. However, as Mankovich et al. (2016) and Mankovich & Fortney (2020) also discuss, Morales temperatures were only calculated for a fraction of helium of . Given that helium is expected to redistribute inside any gas giant planet, a faithful model of the interiors of these planets that takes into account helium rain must be a function of helium abundance. In the following section, we discuss the work of Schöttler & Redmer (2018), which used ab initio methods, including non-ideal effects, but also as a function of helium abundance.

### v2

@@ Abstract
The equation of state of hydrogen-helium (H-He) mixtures plays a vital role in the evolution and structure of gas giant planets and exoplanets. Recent equations of state that account for hydrogen-helium interactions, coupled with hydrogen-helium immiscibility curves, can now produce more physical evolutionary models, such as accounting for helium rain with greater fidelity than in the past. In this work, we present a set of tools for planetary evolution111All tables of thermodynamic quantities and derivatives are available at https://github.com/Rob685/hhe_eos_misc, along with a unified Python interface. Tutorials demonstrating the interface are also available in the repository. that provides a Python interface for existing tables of useful thermodynamic quantities, state-of-the-art H-He equations of state, and pressure-dependent H-He immiscibility curves. In particular, for a collection of independent variable choices, we provide scripts to calculate the variety of thermodynamic derivatives used to model convection and energy transport. These include the chemical potential derived from the internal energy, which is a modeling necessity in the presence of composition gradients when entropy is the other primary variable. Finally, an entropy-based convection formalism is presented and fully described that highlights the physical differences between adiabatic and isentropic interior models. This centralized resource is meant to facilitate both giant planet structural and evolutionary modeling and the entry of new research groups into the field of giant planet modeling.
@@ 1 Introduction
The structure and evolution of gas giant planets and brown dwarfs depend on the properties of hydrogen and helium at high pressures (Hubbard, 1970; Hubbard & Slattery, 1971; Hubbard & Dewitt, 1985; Burrows et al., 1997, 2001; Fortney & Hubbard, 2003, 2004; Militzer et al., 2008; Nettelmann et al., 2008, 2012, 2015; Hubbard & Militzer, 2016; Mankovich et al., 2016; Püstow et al., 2016; Miguel et al., 2016; Debras & Chabrier, 2019; Mankovich & Fortney, 2020; Debras et al., 2021a; Mankovich & Fuller, 2021; Howard et al., 2023), as well as their atmospheres (Burrows et al., 1997; Fortney et al., 2011; Burrows, 2014; Chen et al., 2023). The equation of state (EOS) most employed in the past was published by Saumon et al. (1995, SCvH95). This EOS incorporated the so-called “chemical picture” of the constituent species, obtained by minimizing the Helmholtz free energy of its constituents. This model estimated the gradual dissociation and ionization of hydrogen with increasing pressure, thus accounting for the increasing densities and even the predicted hydrogen first-order phase transition within Jupiter. Subsequent ab initio calculations by Militzer & Hubbard (2013, MH13), who included non-ideal interaction effects of hydrogen-helium mixtures and partial ionizations, were an important step forward. Models using updated EOSes by MH13 and Nettelmann et al. (2012) resulted in significant differences in the interior structures of Jupiter and Saturn (see Miguel et al., 2016; Militzer & Hubbard, 2013; Hubbard & Militzer, 2016; Miguel et al., 2022). Further complications arose when applying these EOSes to such modeling due to the phase separation of helium related to the metallization of hydrogen at high pressures; the presence of a higher-Z core; and the presence of a component of heavy elements in the gaseous mantle. The distribution of the latter is the subject of intense debate following the recent interpretations (Leconte & Chabrier, 2012, 2013; Vazan et al., 2016; Wahl et al., 2017; Folkner et al., 2017; Vazan et al., 2018; Stevenson, 2020; Helled et al., 2022; Militzer et al., 2022; Howard et al., 2023) of the Juno (Bolton et al., 2017a, b) gravity data. To this end, two primary areas of focus are an improved H-He EOS and a more accurate H-He phase separation curve.
@@
The first area, the H-He EOS, demands ab initio calculations of relevant thermodynamic quantities. The computational costs have impeded the creation of a comprehensive tabular EOS for a wide range of hydrogen-helium mixtures, temperatures, and pressures. As a result, current EOSes designed to be employed in structural and evolutionary codes, such as those of SCvH95, Nettelmann et al. (2012); Becker et al. (2014), Chabrier et al. (2019, CMS19) and Mazevet et al. (2022, MLS22), rely on idealized mixing laws to compute the entropy of H-He mixtures. A new class of models includes non-ideal effects of the entropy of mixing (Militzer & Hubbard, 2013; Chabrier & Debras, 2021). These terms have recently been estimated by Howard & Guillot (2023, HG23) based on the calculations of MH13 and CD21, where MH13 calculated a H-He mixture EOS for a helium mass fraction () of 0.245. This resulted in estimates for the non-ideal corrections of the entropy and specific volume (density) for any helium mass fraction for the CMS19 and MLS22 EOSes. However, Howard & Guillot (2023) did not provide the corresponding correction for the internal energy.
@@
Between the different EOS and phase separation calculations by various authors, planetary modelers have a wide range of options. In this paper, we collect many major EOS and miscibility databases required for giant planet evolutionary and structural calculations and describe their strengths and weaknesses222We omit the H-He EOS of Nettelmann et al. (2012) and Becker et al. (2014) due to the lack of entropy calculations available in their tables. An integration technique to obtain the entropy for these tables depends on choices of initial thermodynamic values, making the comparisons between EOSes uncertain and subsequent derivative calculations subject to this uncertainty.. We invert each EOS on a dense grid and calculate all the necessary quantities and derivatives that play a role in the physics, structure, thermal evolution, and convection criteria within Jupiter and Saturn in particular, and giant exoplanets in general. Each EOS is calculated along pressure-temperature, density-temperature, and entropy-density axes that span the thermodynamic range of the interiors of gas giants. As discussed in Section 6, a proper treatment of convection requires the calculation of various derivatives. In addition to the various EOS tables, we furnish useful tables of the hydrogen-helium immiscibility curves of Lorenzen et al. (2009, 2011) and Schöttler & Redmer (2018). While these data are published and employed in some of the most recent planetary evolution calculations for Jupiter and Saturn (Nettelmann et al., 2015; Püstow et al., 2016; Mankovich et al., 2016; Mankovich & Fortney, 2020), smooth fits to these data are not easily come by. Though demixing temperatures for a broad range of helium abundances are what authors such as Lorenzen et al. (2009, 2011) and Schöttler & Redmer (2018) have provided, easily used fits to these data are not easily generated. We provide such smooth fits here. In this way, our work provides a more centralized and standardized approach by collecting the most recent EOS and H-He immiscibility data in a self-consistent format.
This paper is organized as follows. Section 2 discusses the various hydrogen-helium (H-He) EOSes, compares their properties, and discusses their applications. In Section 3, we provide an overview of the available miscibility calculations in the literature, compare the different demixing temperature calculations of each of the sources, compare the H-He demixing temperatures with different available H-He equations of state, and discuss what various authors have concluded concerning helium rain in Jupiter and Saturn based on these different demixing temperatures. Section 4 describes the different EOS table inversions, inversion techniques, interpolation methods, and thermodynamic table ranges. This section also describes the H-He EOS and miscibility code structure. In Section 5, we briefly discuss the thermodynamic consistency of the EOSes published by Saumon et al. (1995), citeChabrier2021, and Chabrier et al. (2019) (with the non-ideal corrections by Howard & Guillot, 2023)). In Section 6, we discuss the various derivatives necessary for convection and energy calculations in a planetary model. Section 7 showcases an example evolutionary calculation using the code APPLE (Sur et al., 2024) and the tables presented here. Finally, in Section 8, we provide a summary discussion and concluding remarks. In the appendices, we derive various of the relationships provided in the text.
@@ 2 Hydrogen and Helium Equations of State
Historically, the EOS of hydrogen and helium at high pressures has been divided into the “physical” picture and the “chemical” picture. In the physical picture, electrons and nuclei interact through Coulomb potentials, and the EOS is obtained by solving the resulting quantum many-body problem. In the chemical picture, atoms remain bound and interact through pair potentials. However, at ionization pressures and temperatures, electrons become unbound and delocalized. In the high temperature regime, electrons become unbound and the physical picture is tractable (Rogers, 1981; Rogers & Young, 1984). However, in the low-temperature and high-pressure regime relevant to gas giant planets, the physical picture could not easily handle partial ionization. The chemical picture presents a simplified model a physical model. In the past, the chemical picture of SCvH95 was the most commonly used H-He EOS. However, with improvements in computation rigorous physical (or “ab initio”) models have now been published. Most notably, MH13, CMS19, Chabrier & Debras (2021, CD21), and MLS22 have published EOSes obtained from ab initio calculations. All of these models rely on the additive volume law (AVL), since calculations for arbitrary hydrogen-helium mixtures remain intractable. This approximation was improved by HG23, who have calculated nonideal interaction corrections based on the MH13 results for the CMS19 and MLS22 EOSes. In this section, we discuss the properties of the historical SCvH95, the ab initio MH13, and the most c urrent CMS19 EOSes and compare their properties333The MLS22 EOS result is a hydrogen EOS table benchmarked with structural calculations of the interior of Jupiter inferred from the Juno data. Though we present this EOS and provide the same quantities and derivatives for it, we focus here on the CMS19 and SCvH95 EOSes..
@@ 2.1 Saumon, Chabrier, and van Horn (1995)
Saumon et al. (1995, SCvH95) employed the chemical picture and generated an EOS that covered a temperature range of and a density range of . They calculated pure hydrogen and helium tables and relied on the AVL, to handle a general helium mass fraction dependence. Intensive variables, such as temperature and pressure, must be uniform throughout a system at equilibrium. Extensive variables, such as volume and entropy, are additive for a system in equilibrium. Therefore, the AVL considers a combined system of equations of state and assumes they are in thermal and pressure equilibrium. AVL cannot account for interactions between hydrogen and helium, is not reliable in regions of partial ionization, and it does not account for phase separation (see Section 6 of SCvH95 for more detailed descriptions of its limitations). Despite these limitations, until recently, SCvH95 was the most used EOS in the study of low-mass stars, brown dwarfs, and giant planets (e.g., Burrows et al., 1995; Guillot et al., 1996; Saumon et al., 1996; Burrows et al., 1997; Chabrier & Baraffe, 1997; Baraffe et al., 1998; Burrows et al., 2001; Fortney & Hubbard, 2003, 2004; Fortney et al., 2011)
@@ 2.2 Militzer & Hubbard (2013)
The Militzer & Hubbard (2013, MH13) EOS spans a temperature range of K and a density range of g cm-3. As their model is a physical model, they avoid the limitations of the low-temperature and high-pressure regime inherent in the chemical model of SCvH95. They use Density Functional Theory Molecular Dynamics (DFT-MD), coupled with the thermodynamic integration (TDI) technique, to obtain the entropy directly through the difference in Helmholtz free energy between hydrogen and helium mixtures while using the Perdew-Burke-Ernzerhof (Perdew et al., 1996, PBE) exchange-correlation functional. To account for interactions between species and electrons via TDI, they introduced classical intramolecular pair potentials to match their DFT calculations as closely as possible (Militzer & Hubbard, 2013).
@@ 2.3 Chabrier, Mazevet, and Soubiran (2019)
Chabrier et al. (2019, CMS19) constructed pure hydrogen and helium EOS tables spanning wider density and temperature ranges, namely g cm-3 and K. For densities lower than 0.05 g cm-3, they use the SCvH95 hydrogen EOS, the hydrogen EOS of Caillabet et al. (2011) for g cm-3, and the EOS of Chabrier & Potekhin (1998) for densities greater than 10 g cm-3.
As does SCvH95, their EOS relies on the AVL for helium mixtures and includes only the ideal entropy of mixing. CD21 incorporate the non-ideal interaction effects of MH13 into the CMS19 EOS in a manner similar to Miguel et al. (2016). Whereas Miguel et al. (2016) did not account for the non-ideal entropy contributions from MH13, CD21 did. Using the AVL, CD21 obtained an “effective” hydrogen EOS from MH13 using the CMS19 helium EOS to obtain a mixture EOS that agrees with MH13 at (see Figure 9 of CD21).
@@
and requiring that these corrections agree with MH13 at and with CD21 at . The corrections, and , are assumed to be proportional to the products of hydrogen and helium mass fractions.
The CMS19 EOS with the HG23 corrections (CMS19+HG23) and the CD21 EOS are the most current development of an H-He EOS that spans all relevant thermodynamic space, helium fractions, and nonideal interactions. In Figure 1, we compare the CMS19+HG13, CD21, and MLS22+HG23 EOSes with the SCvH95, MH13, and CMS19 tables, illustrating the differences among them. The top panel mirrors 2 of Sur et al. (2024), originally inspired by HG23, and compares the EOSes at a helium fraction of and a specific entropy of baryon. The bottom panel shows the temperature differences due to the non-ideal interactions at different entropy values, as noted above, for each profile. The density deviations from SCvH95 increase with decreasing entropy, particularly between 0.5 and 7 Mbar. This pressure region is expected to exhibit H-He phase separation at a sufficiently low temperature (Stevenson, 1975; Stevenson & Salpeter, 1977a). Figure 1 also compares the EOS adiabats to a 1 MJ profile at 0.5 and 5 Gyrs using APPLE (Sur et al., 2024) with the CD21 EOS. A full example of a 1 MJ evolution with APPLE is discussed in Section 7. In the next section, we discuss the current understanding of H-He immiscibility in this context.
@@ 3 Hydrogen and Helium Immiscibility
Early evolutionary calculations for Saturn (Pollack & Graboske, 1976; Pollack et al., 1977) could not reproduce its current heat flux. To account for this apparent excess heat, Stevenson (1975) proposed the formation of He-rich droplets that rain into the interior (also see Stevenson & Salpeter, 1977a). One of the first attempts to capture and apply the demixing of hydrogen and helium as a function of helium abundance was done by Hubbard & Dewitt (1985), who applied Monte Carlo simulations for a fully ionized hydrogen and helium plasma. The low-pressure regime was studied by Schouten et al. (1991) via Monte Carlo methods using pair potentials for interactions. Pfaffenzeller et al. (1995) calculated the demixing temperatures using molecular dynamics (MD) and found a demixing temperature-pressure dependence opposite that of Hubbard & Dewitt (1985). However, the demixing temperatures remained too cold for Saturn to ever reach the immiscibility regions (Pfaffenzeller et al., 1995; Fortney & Hubbard, 2003). Fortney & Hubbard (2003) created an ad hoc phase diagram, adjusting the demixing temperatures of Pfaffenzeller et al. (1995), to match the present effective temperature of Saturn. More recently, H-He phase diagrams have been calculated by Lorenzen et al. (2009, 2011, L0911), Morales et al. (2009, 2013), and Schöttler & Redmer (2018, SR18). These ab initio phase diagrams are discussed in the next sections.
@@
Lorenzen et al. (2009, L0911) conducted the first ab initio immiscibility gap model. They used DFT-MD with a PBE exchange-correlation functional and pair potentials for hydrogen and helium. They calculated the demixing temperatures for pressures between 4 Mbar and 24 Mbar and expanded their model in Lorenzen et al. (2011) to include 1 and 2 Mbar. In all cases, they relied on linear mixing to calculate the enthalpy of mixing to then infer the change in the Gibbs free energy, . This was then minimized to constrain the H-He miscibility gaps. They linearized the change in the enthalpy with
@@
where ] is the change in the ideal entropy of mixing.
@@
It has been suggested by both Püstow et al. (2016) and Mankovich et al. (2016) that the unperturbed phase diagrams of L0911 are unlikely due to their high demixing temperatures. Püstow et al. (2016) found that to obtain for Saturn an atmospheric helium mass fraction between and a shift of the L0911 miscibility curve needed to range from K to K, respectively. For Jupiter, Mankovich et al. (2016) found that to fit the observed value of (Von Zahn et al., 1998) this median shift of K. In the next section, we discuss the work of Morales et al. (2009, 2013) and their inclusion of non-ideal interactions in the phase diagram.
@@
Morales et al. (2009) constructed the first H-He phase diagram that included non-ideal interaction effects in the entropy of mixing, following a similar methodology to that of Militzer (2009) and MH13. Morales et al. (2009) used the Born-Oppenheimer MD coupled with the PBE exchange-correlation functional, and used TDI to infer the entropy using the Helmholtz free energy difference (also used by MH13). Morales et al. (2013) expanded their work to pressures below 0.4 Mbar, and claimed that the linear mixing assumption leads to significant disagreements with their direct calculation methods at high helium mass fractions and high pressures.
This work led to colder demixing temperatures compared to L0911. While Jupiter had a small miscibility gap between Mbar and Mbar with the L0911 phase diagram, the Morales demixing temperature results were too cold to allow Jupiter any miscibility gap (Morales et al., 2013). For consistency, Hubbard & Militzer (2016) used the Morales phase diagram, since both the EOS and the demixing temperatures were calculated in a similar fashion. Hubbard & Militzer (2016) then claimed that helium rain occurs within Jupiter at an entropy value of baryon. However, as Mankovich & Fortney (2020) also discuss, the published demixing temperatures are restricted to only the solar helium number fraction of . Given that helium is expected to redistribute inside any gas giant planet, a faithful model of the interiors of these planets that takes into account helium rain must be a function of helium abundance. In the following section, we discuss the work of Schöttler & Redmer (2018), which used ab initio methods, including non-ideal effects, but also as a function of helium abundance.